OpenStudy (anonymous):

fx=1+(1/x)-(1/x^2) f'=0+(1/1)-(1/1^2) or 0? Please help..

OpenStudy (anonymous):

The question is unclear. Are you supposed to find the derivative of \[f(x)=1+\frac{1}{x}-\frac{1}{x^2}\] and evaluate it for \(x=1\) or something?

OpenStudy (anonymous):

Yes, in order to set x=0 for critical points.

OpenStudy (anonymous):

Well the derivative would be \[f'(x)=-\frac{1}{x^2}+\frac{2}{x^3}\] If you want to find the critical points, set it to 0 and solve for \(x\): \[\frac{2}{x^3}-\frac{1}{x^2}=0~~\iff~~x^3\left(\frac{2}{x^3}-\frac{1}{x^2}\right)=0\cdot x^3~~\iff~~2-x=0\]

OpenStudy (anonymous):

Thank you, let me work on this. Thanks again.

OpenStudy (anonymous):

yw